find joint equation of line passsing through origin and one is parallel to line3x+2y+5 is equal to zer


and other is perpendicular to 5x+3y-2 is equal to zero

Grade:

1 Answers

Chetan Mandayam Nayakar
11 years ago

first line: slope=-3/2, thus equation is y=-3x/2, y+(3x/2)=0

second line: slope=3/5,equation is y=3x/5, y-(3x/5)=0

joint equation is (y+(3x/2))(y-(3x/5))=0